OpenStudy (anonymous):

classify the equations 2x + y - 3=0 x + y - 1 = 0 parallel coincident intersecting

OpenStudy (texaschic101):

2x + y - 3 = 0 y = -2x + 3 x + y - 1 = 0 y = -x + 1 different slopes (so its not parallel or coincident), therefore, it is intersecting

OpenStudy (anonymous):

Thank You So Much !

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

Can You Please Help Me On The Next Ones?